find slope of a line perpendicular to the line that passes through points
(3,9) and (7, 15)

Answer by dfrazzetto(283)

First find the slope of the original line
slope = m1 = y2-y1/(x2-x1) = 15-9/(7-3) = 6/4 = 3/2
A line of perpendicular slope will have slope m2 such that it is the negative reciprocal of the original slope m1
m2 = inverse reciprocal of 3/3 = -2/3